OpenStudy (anonymous):

kyle is buying a new car. At Bank A, a car loan will cost him 11.5%. At bank B, it will cost 12.25%. if he borrows $8,000, how much will he save if he borrows from bank A ?

OpenStudy (texaschic101):

Bank A : (8000)(.115) = 920 -- this is your interest 8000 + 920 = 8920 -- this is what has to be paid back Bank B : (8000)(.1225) = 980 -- this is your interest 8000 + 980 =8,980 -- this is what has to be paid back 8980 - 8920 = 60 ANSWER : Kyle will save $ 60 if he borrows from Bank A.

OpenStudy (anonymous):

that was right (: !! your good at explaining

OpenStudy (texaschic101):

thank you....I try....I have to go now....I wish you the best of luck in your classes. By the way....you have a cool username :) Have a nice night

OpenStudy (anonymous):

okayy thanks (: and talk to you soon !

OpenStudy (texaschic101):

ok...bye